# SpikeLLM: Scaling up Spiking Neural Network to Large Language Models via Saliency-based Spiking

### v1

@@ 4 Spike-Driven Quantization
Towards accurate quantized LLMs, we conquer outlier and salient values via an Optimal Brain Spiking framework. Before that, we first define a Generalized Integrate-and-Fire (GIF) neuron as an alternative to a traditional quantizer, which is equivalent to inferring in binary event-driven or low-bit workloads. SNNs with 7~70B parameters are efficiently built up based on Optimal Brain Spiking.
@@ 4.1 Generalized Integrate-and-Fire Neuron
As shown in Eq.3, 4, the original IF neuron encodes one binary spike each step. To fully express FP16 values by the firing rate encoding, $2^{16}$ spiking steps are required, which is much more inefficient compared with 16-bit quantization. Based on the inefficiency of IF spike encoding, recent SNNs are almost less than 1 billion parameters and are hard to train with long-time backpropagation through time (BPTT). On the other hand, traditional quantization encodes all binary digits in one step, which makes it hard to quantize outliers. Based on both two aspects, we make a balance between auto-regressive steps and encoding length in each step. This is achieved by merging $L$ spiking steps and encoding each step as low-bit digits with $log_{2}L$ length. We define this L-step merged IF neuron as the Generalized Integrate-and-Fire (GIF) neuron:
$\mathbf{s}_{GIF}(t)=\frac{1}{L}\sum_{t=1}^{L}\mathbf{s}_{IF}(t),\>\mathbf{s}_{GIF}(t)=\begin{cases}\frac{k}{L},&\text{ if }kV_{th}\leq L\mathbf{v}(t)<(k+1)V_{th},k=0,1,...,L-1\\ 1,&\text{ if }\mathbf{v}(t)\geq V_{th}\end{cases}.$ (7)
In each spiking step, there are L quantization levels in the low-bit workloads, while GIF can also expand back to binary workloads as IF to become event-driven. In general, for larger merging-step $L$, low-bit workloads are more efficient when measured by ACE [58], while for smaller, binary workloads are more efficient caused by event-driven sparsity.
Remark 1. Ternary Spike. The ternary spike $\mathbf{s}_{Ter}(t)$ proposed by SpikeLM [52] is a special case of GIF, which is formulated by merging a positive IF neuron $\mathbf{s}^{+}_{IF}(t)$ and a negative $\mathbf{s}^{-}_{IF}(t)$. Ternary spike not only increases quantization levels but also keeps additive in SNNs.
@@
Figure 1: Per-channel spiking mechanisms in SpikeLLM. (Left) Per-channel spiking self-attention. Salient channels in the KV caches are quantized by multi-step spikes, while the value and attention map copy part channels accordingly. (Right) Linears with per-channel spiking activations or weights. We detect salient channels by gradient or hessian metric for activations or weights respectively.
@@ 4.2 Optimal Brain Spiking
Traditional quantization leverages per-channel smooth techniques for activation outliers and mix-precision quantization for salient weights respectively. Thanks to the auto-regressive encoding in spiking neurons, we address accurate quantization for both salient activations and weights in a unified algorithm. As shown in Fig. 1, we apply GIF neurons as quantizers for self-attentions (including the query and key-value caches), activations, and weights.
Based on the fact that not all channels are equally important in LLMs, a per-channel spiking mechanism is proposed to address outlier and salient value quantization issues in Section 3.2. This is addressed by expanding salient channels in activations or weights with more spiking steps compared with unimportant channels. Given the GIF neuron to quantize activations, we first detect salient channels $C$ and the other channels $C^{\prime}$, and then allocate $T$-step spikes to simulate channels in $C$ and fewer $T^{\prime}$-step (usually $T^{\prime}=1$) for others in $C^{\prime}$, and the spiking linear layer can be represented as:
${\mathbf{x}}^{(\ell)}$ $=\frac{V_{th}}{T}\sum_{t=1}^{T}\mathbf{W}\mathbf{s}^{(\ell)}(t)+{\mathbf{W}\min(\mathbf{x}^{(\ell)})}+\mathbf{b}$ (9)
$\simeq\frac{V_{th}}{T}\sum_{t=1}^{T}\mathbf{W}\mathbf{s}^{(\ell)}(t)|_{\mathbf{s}\in C}+\frac{V^{\prime}_{th}}{T^{\prime}}\sum_{t=1}^{T^{\prime}}\mathbf{W}\mathbf{s}^{(\ell)}(t)V_{th}^{\prime}|_{\mathbf{s}\in C^{\prime}}+{\mathbf{W}\min(\mathbf{x}^{(\ell)})}+\mathbf{b},$
where $V_{th}=\frac{\max{(\mathbf{x}_{C})}}{T}$ and $V_{th}^{\prime}=\frac{\max{(\mathbf{x}_{C^{\prime}})}}{T^{\prime}}$ are per-channel spiking thresholds in $C$ and $C^{\prime}$, which confirms not clipping max values (shown in Eq.6). As shown in Fig. 1, we apply this per-channel spiking mechanism in KV-caches and activations, or weights. If the salient channels in KV-caches and activations spiking T steps, the other side of the matrix multiplication keeps one-step quantization, and copies part of channels for the same T steps accordingly. In turn, if the salient channels in weights spiking, the corresponding channels in activations are copied. Thus, it is essential to detect the salient channels in both weights and activations, and an Optimal Brain Spiking framework is proposed.
Our Optimal Brain Spiking is a weight-activation generalization of the classic Optimal Brain Surgeon (OBS) framework [21]. Different from OBS which focuses on weight pruning with only the second-order differentiation, we focus on detecting salient channels in both activations and weights via both first and second-order differentiation. Given a post-training model well-optimized under a loss function $\mathcal{L}$, any weights or activations $\mathbf{x}$ in the model can be expressed by a second-order Taylor expansion around its optimal value $\mathbf{x}^{*}$:
@@
For weights, the gradient is zero because the optimizer directly optimizes weights to the local minimum after pretraining. Thus, the first-order term is zero in Eq.10 and $\delta\mathcal{L}(\delta\mathbf{w})$ has to approximate via the second-order term in Taylor expansion: $\delta\mathcal{L}(\delta\mathbf{w})\simeq\frac{1}{2}\delta\mathbf{w}^{\top}\mathbf{H_{\mathcal{L}}}(\mathbf{w}^{*})\delta\mathbf{w}$, which is proved in OBS [21]. And we apply the same weight saliency metric as OBS-based methods [46, 16, 14]:
$\text{Saliency}(\mathbf{W}_{ij})=\frac{\mathbf{W}_{ij}^{2}}{[\mathbf{H}_{ii}^{-1}]^{2}},\quad\mathbf{H}=\mathbf{X}\mathbf{X}^{\top}.$ (12)
@@
The saliency matrix $\text{Saliency}(\mathbf{X})$ and $\text{Saliency}(\mathbf{W})$ have the same shape with activations $\mathbf{X}$ and weights $\mathbf{W}$, which are inefficient to store. As shown in Fig.2, we calculate per-channel or per-token means of the saliency matrix and find both the per-channel saliency in activations and weights is robust enough to detect salient channels while per-token is insignificant. Based on Eq.11, 12 with a few calibration data, we first compute per-channel saliency and generate masks to select the most salient channels in Eq.9, and then store these lightweight masks for inference.
Figure 2: Comparisons of first- (or second-) order saliency metrics in the first linear layer. (Left) For activations, the first-order gradient saliency is insignificant in the token dimension. (Mid) For activations, the first-order gradient saliency is significant in the channel dimension. (Right) For weights, the second-order Hessian saliency is significant in the channel dimension.
@@ 5 Experiments
We evaluate both the effectiveness and efficiency of the brain-scale SpikeLLM with 7~70 B parameters from two aspects: (i) general weight-activation quantization in very low-bits is our main focus; (ii) towards additive large language models, binary or ternary quantization is also performed.
Meraging Steps and Spiking Steps. To simulate M-bit quantization, we set $L=2^{M}$ in the GIF neuron 7 to make it have the same quantization levels. For salient channels, GIF neurons auto-regressively encode for T steps, and each step has L levels.
Training Details. We follow main streams of post-training quantization (PTQ) [16] and calibration [47] pipelines. (i) For low-bit quantization setting, our primary baseline in OmniQuant [47] and we keep the same training settings. In details, we randomly select 128 calibration data from WikiText2, which are 2048-token chunks. We select LLAMA-1 (7B) [48] and LLAMA-2 (7B, 13B, 70B) [49] as full-precision base models. The layerwise calibration is used and 4W4A (4-bit activation, 4-bit weight), 2W8A, or 2W16A quantizations are trained. We report full details pipelines in Appendix A. (ii) For the addition SpikeLLM setting, we follow the GPTQ [16] framework using the same 128 WikeText2 calibration data. With the same setting as PB-LLM, we use LLAMA-2-7B and set a similar cost.
Evaluation Tasks. Following our primary baselines, OmniQuant [47] and PB-LLM [46], we evaluate preplexity (PPL) of language generation in WikiText2 [37] and C4 [42] benchmarks. As in previous works, we report zero-shot accuracy in PIQA [3], ARC-easy [8], ARC-challenge [8], BoolQ [7], HellaSwag [8], and Winogrande [44]. The same evaluation methods with OmniQuant are applied.
@@
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 0.0687$\times$ 65.83 47.98 27.82 64.22 43.49 53.28 50.44
LLAMA-2-7B FP16 1$\times$ 78.45 69.32 40.02 71.07 56.69 67.25 63.80
OmniQuant W4A4 0.0625$\times$ 62.19 45.62 25.43 60.89 39.15 52.17 47.58
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 0.0687$\times$ 64.47 48.74 27.30 63.27 43.29 56.83 50.65
OmniQuant W2A8 0.0625$\times$ 51.90 27.82 19.97 38.26 25.85 50.36 35.69
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A8 0.075$\times$ 54.62 30.89 20.90 53.91 30.75 53.12 40.70
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 0.138$\times$ 65.61 48.15 27.39 60.46 39.01 52.80 48.90
LLAMA-2-13B FP16 1$\times$ 78.78 73.36 45.56 68.99 59.73 69.61 66.01
OmniQuant W4A4 0.0625$\times$ 67.03 53.96 30.55 62.91 44.83 53.91 52.20
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 0.0687$\times$ 66.49 55.30 30.12 64.16 47.43 51.54 52.49
OmniQuant W2A8 0.0625$\times$ 57.51 35.27 19.11 61.31 29.95 49.41 42.09
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A8 0.075$\times$ 64.09 48.74 24.23 62.29 40.38 52.49 48.70
OmniQuant W2A16 0.125$\times$ 63.55 45.16 23.55 62.45 39.84 53.12 47.95
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 0.138$\times$ 67.63 53.70 28.33 63.64 45.10 57.22 52.60
LLAMA-2-70B FP16 1$\times$ 81.07 77.74 51.11 76.70 63.99 77.03 71.27
OmniQuant W2A16 0.125$\times$ 62.57 43.86 22.78 56.42 39.60 52.49 46.29
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 0.138$\times$ 76.44 66.92 38.31 66.88 51.86 59.19 59.93
Table 2: Comparisons between SpikeLLM and OmniQuant in the same pipeline with the Wikitext2 and C4 PPL metrics. We do not evaluate the W4A4 and W2A8 settings for LLAMA2-70B because the grouped-query attention (GQA) makes training unstable in the OmniQuant pipeline.
LLAMA-2-7B LLAMA-2-13B LLAMA-2-70B
Method Saliency #Bits Wikitext2 C4 Wikitext2 C4 Wikitext2 C4
OmniQuant W4A4 15.25 19.35 12.40 15.87
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 11.36 15.87 9.71 12.10
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W4A4 11.41 14.34 9.75 12.17
OmniQuant W2A8 287.64 445.21 53.87 72.33
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A8 22.13 30.45 13.56 18.73
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A8 28.78 44.80 12.80 17.05
OmniQuant W2A16 38.05 98.74 17.14 27.12 10.04 19.31
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 14.16 19.73 9.45 13.86 6.35 9.62
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A16 14.50 19.82 9.17 12.37
5.1 Low-Bit Quantization Results
As shown in Table 1, we compare SpikeLLM with state-of-the-art weight-activation quantization methods including SmoothQuant, LLM-QAT, and OmniQuant, which shows SpikeLLM improves significantly based on OmniQuant with a few additional spikes to enhance salient channels. For LLAMA-2-7/13B in Table 1, this performance enhancement is dramatic: for example, for the 2W8A and 2W16A quantization of 7B, improvements are 5.01% and 7.82% respectively with 20% and 10% additional spikes. In Table 2, their WikiText2 PPL of $\text{SpikeLLM}_{\texttt{T=2}}$ also significantly decrease 92.31% and 62.79% compared with baselines.
Next, we confirm the efficiency of SpikeLLM with almost the same operations compared with quantization. As shown in Fig.3 (Left), We increase average spiking steps in activations and decrease in weights to keep almost the same operations, which indicates SpikeLLM exceeds all of its equal-operation settings. To evaluate the efficiency of the Optimal Brain Spiking framework, as shown in Fig.3 (Mid, Right), we compare SpikeLLM with equal-operation baselines with randomly selected additional spiking channels in weights and activations respectively, which show stable better performance for SpikeLLM. The first- or second-order differentiation-based saliency methods are crucial for activation or weight quantization. Based on these observations, we report our proposed Optimal Brain Spiking framework (as well as GIF neurons) is accurate and robust enough to enhance or even take place traditional quantization methods for both weight and activation quantization.
Figure 3: Effiency of Generalized Integrate-and-Fire neurons and Optimal Brain Spiking in LLAMA-2-7B. (Left) We keep the same computational cost in different GIF settings and compare it with the OmniQuant-4A4W baseline (in red). (Mid) We compare Optimal Brain Spiking with randomly increased spiking steps in activations with the same cost. SpikeLLM or SpikeLLM-KV indicates two settings: spiking activations or spiking both activations and KV-caches. We do not randomly increase spiking steps in KV-caches for the Random-Spike-4A4W baseline, because it is unstable and occurs N/A gradient. (Right) We compare Optimal Brain Spiking with randomly increased spiking steps in weights. SpikeLLM indicates increasing spiking steps as 2 in parts of weight channels.

### v2

@@ 4 Spike-Driven Quantization
To avoid inefficient spike encoding in previous SNNs and significant quantization error in Quantized ANNs, we propose the first spiking large language model with two techniques: Generalized Integrate-and-Fire neurons for spike length comparison and Optimal Brain Spiking framework for accurate saliency-based spike encoding. Compared with IF-SNNs, SpikeLLM compresses binary code length from T to approximate $log_{2}T$ and can infer with both the low-bit matrix multiplication workload on GPUs and the per-bit computation workload on neuromorphic chips.
@@ 4.1 Generalized Integrate-and-Fire Neuron
Because of the inefficiency of IF neurons, recent SNNs are almost less than 1 billion parameters and are hard to train with long-time backpropagation through time (BPTT). On the other hand, traditional quantization encodes all binary digits in one step, which makes it hard to quantize outliers. Based on both aspects, we make a balance between recursive steps and code length in each step. This is achieved by merging $L$ spiking steps and encoding each step as low-bit digits with $log_{2}L$ bit length. We define this L-step merged IF neuron as the Generalized Integrate-and-Fire (GIF) neuron:
$\mathbf{s}_{GIF}(t^{\prime})=\sum_{t=1}^{L}\mathbf{s}_{IF}(t),\>\mathbf{s}_{GIF}(t^{\prime})=\begin{cases}k,&\text{ if }kV_{th}\leq L\mathbf{v}(t^{\prime})<(k+1)V_{th},k=0,1,...,L-1\\ L,&\text{ if }\mathbf{v}(t)\geq V_{th}\end{cases},$ (5)
where there are L quantization levels in each spiking step. After merging, the recursive steps $T^{\prime}=\frac{T}{L}$, $t^{\prime}=\lfloor\frac{t}{L}\rfloor$ and each step has been encoded by $log_{2}L$ bit quantization.
According to Eq.5, if simulating 32 ($T=32$) quantization levels (or spiking steps), we can set the merged spiking step $T^{\prime}=2$ and spiking level $L=16$ in each step. After merging, each step is quantized to $4\times\{0,1\}$ digits, and the total 32 quantization levels can be represented by $8\times\{0,1\}$ digits. Although GIF neuron still has longer code than traditional quantization, this gives us the freedom to choose spiking steps according to channel saliency in LLMs and achieve higher performance with similar costs.
Remark 1. Ternary Spike. The ternary spike $\mathbf{s}_{Ter}(t)$ proposed by SpikeLM (Xing et al., 2024b) is a special case of GIF, which is formulated by merging a positive IF neuron $\mathbf{s}^{+}_{IF}(t)$ and a negative $\mathbf{s}^{-}_{IF}(t)$. Ternary spike not only increases quantization levels but also keeps additive in SNNs.
@@
On-Chip Inference. Due to the equivalence before and after spike merging in Eq. 5, we can choose different workflows for training and inference: (i) both multi-bit training and inference on GPUs: as shown in Figure 1 (d), GIF neurons can be converted to single-precision quantization and apply the general matrix multiplication kernels for inference, which equals to quantization but improves performance in Table 1. (ii) Multi-bit training on GPUs and 1-bit inference on neuromorphic chips: during training, merged spikes are employed, while during inference, the merged spikes are equivalent to expand back into their 1-bit formulation before merging. For example, if the value is $n$ in a step, it can be decoded as $n$ spikes occurred in $L$ time. This workflow avoids long-distance BPTT during training and therefore enhances training accuracy. We detial the training and inference of GIF neurons in Appendix A.4.
Figure 2: saliency-Aware spiking mechanisms in SpikeLLM. (Left) Spiking self-attention. Salient channels in the KV caches are encoded by multi-step spikes. (Right) Spiking activations or weights in a linear layer, where saliency is detected by gradient or Hessian metric respectively.
4.2 Saliency-Aware Spiking Steps
Given the unequal importance of channels in LLMs, we propose a saliency-aware spiking mechanism to quantify this concept guiding the subsequent quantization process. This is addressed by expanding salient channels in activations or weights with more spiking steps compared with unimportant channels. Given the GIF neuron to quantize activations, we first detect salient channels $C^{\prime}$ and the other channels $C$, and then allocate $T^{\prime}$-step spikes to encode channels in $C^{\prime}$ and one-step for others in $C$, which can be represented by:
${\mathbf{x}}^{(\ell)}$ $=\frac{V_{th}}{T^{\prime}}\sum_{t=1}^{T^{\prime}}\mathbf{s}^{(\ell)}(t)+{\min(\mathbf{x}^{(\ell-1)})}$ (7)
$\simeq\frac{V_{th}}{T^{\prime}}\sum_{t=1}^{T^{\prime}}\mathbf{s}^{(\ell)}(t)|_{\mathbf{s}\in C^{\prime}}+\mathbf{s}^{(\ell)}(1)V_{th}|_{\mathbf{s}\in C}+{\min(\mathbf{x}^{(\ell-1)})},$
where $V_{th}=\frac{\max{(\mathbf{x})}}{T^{\prime}}$ is per-token spiking thresholds, which confirms not clipping max values (as Eq.4). As shown in Fig. 2, this per-channel spiking mechanism can apply to KV-caches, activations, and weights. For weight-activation quantization, salient channels in KV-caches and activations have $T^{\prime}$ spiking steps, while the other side of the matrix multiplication keeps one-step quantization, and copies corresponding channels for the same $T^{\prime}$ steps. For weight-only quantization, the salient channels in weights have $T^{\prime}$ spiking steps, while the corresponding channels in activations are copied. Following, it is essential to detect the salient channels in both weights and activations.
@@ 4.3 Optimal Brain Spiking
Our Optimal Brain Spiking is a weight-activation generalization of the classic Optimal Brain Surgeon (OBS) framework (Hassibi & Stork, 1992). Different from OBS which focuses on weight pruning with only the second-order differentiation, we focus on detecting salient channels in both activations and weights via both first and second-order differentiation. Given a post-training model well-optimized under a loss function $\mathcal{L}$, any weights or activations $\mathbf{x}$ in the model can be expressed by a second-order Taylor expansion around its optimal value $\mathbf{x}^{*}$:
@@
For weights, the gradient is zero because the optimizer directly optimizes weights to the local minimum after pretraining. Thus, the first-order term is zero in Eq.8 and $\delta\mathcal{L}(\delta\mathbf{w})$ has to approximate via the second-order term in Taylor expansion: $\delta\mathcal{L}(\delta\mathbf{w})\simeq\frac{1}{2}\delta\mathbf{w}^{\top}\mathbf{H_{\mathcal{L}}}(\mathbf{w}^{*})\delta\mathbf{w}$, which is proved in OBS (Hassibi & Stork, 1992). And we apply the same weight saliency metric as OBS-based methods (Shang et al., 2023; Frantar et al., 2022; Frantar & Alistarh, 2022):
$\text{Saliency}(\mathbf{W}_{ij})=\frac{\mathbf{W}_{ij}^{2}}{[\mathbf{H}_{ii}^{-1}]^{2}},\quad\mathbf{H}=2\mathbf{X}\mathbf{X}^{\top}.$ (10)
Figure 3: Comparisons of different saliency metrics in the first linear layer. (a) Insignificant per-token gradient saliency in activations. (b) Significant per-channel gradient saliency in activations. (c) Significant per-channel Hessian saliency in weights. The horizontal axis represents each channel.
@@
In implentation, the saliency matrix $\text{Saliency}(\mathbf{X})$ and $\text{Saliency}(\mathbf{W})$ have the same shape with activations $\mathbf{X}$ and weights $\mathbf{W}$, which are inefficient to store. As shown in Fig.3, we calculate per-channel or per-token means of the saliency matrix. We observe that both the per-channel saliency in activations and weights are robust enough to detect salient channels while per-token is insignificant. Based on Remark 2, we first compute per-channel saliency with calibration data and generate their rank to select the salient channels in Eq.7. Then, we store these lightweight masks for inference.
@@ 5 Experiments
Table 2: Spiking settings to simulate quantization. We set 10% or 5% salient channels for 2 or 4 spiking steps in OBSpiking. Attention, Act. and Weight indicate where to apply OBSpiking.
Quantizer Spike-Level L Step T’ Salient Channels Attention Act. Weight
4bit 16 2 10%
4bit 16 4 5%
2bit 4 2 10%
2bit 4 4 5%
We evaluate the necessity to introduce SpikeLLM due to the inefficiency and incompatibility caused by the direct application of existing ANNs based quantisation strategies to spike neural networks. This necessity is studied and verified in two main directions: (i) general weight-activation quantization in very low bits; (ii) ternary quantized LLM towards additive linear layers.
Training. As shown in Table 2, SpikeLLM can simulate W4A4 (4-bit activation, 4-bit weight), W2A8, or W2A16 quantization with different hyper-parameters, including the spiking step $T^{\prime}$ and spike-level $L$ in GIF neurons, and the ratio of saliency channels in OBSpiking. We follow main streams of post-training quantization (PTQ) (Frantar et al., 2022) and calibration (Shao et al., 2023) pipelines. (i) For the low-bit quantization setting, our primary baseline is OmniQuant (Shao et al., 2023) and we report the detailed pipeline in Appendix A.2. We select LLAMA-1-7B (Touvron et al., 2023a) and LLAMA-2-7B, 13B, 70B (Touvron et al., 2023b) as full-precision models. (ii) For the additive SpikeLLM setting, we follow the GPTQ (Frantar et al., 2022) framework and report training details in Appendix A.3. Both pipelines use the same 128 WikeText2 calibration data for continue layer-wise training based on full-precision LLMs. It takes around 3.5 hours to train a 7B model with OmniQuant pipeline on a single A100-80 GPU.
Evaluation. We follow the same evaluation methods as the primary baselines, OmniQuant (Shao et al., 2023) and PB-LLM (Shang et al., 2023). We evaluate the perplexity (PPL) of language generation in WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) benchmarks. We also evaluate zero-shot common scene reasoning tasks including PIQA (Bisk et al., 2020), ARC-easy (Clark et al., 2018), ARC-challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Clark et al., 2018), and Winogrande (Sakaguchi et al., 2021) datasets.
Operation Metric. (i) For GEMM kernels, we make a direct comparison of SpikeLLM and quantized ANNs by introducing the ACE metric (Zhang et al., 2022b) (Appendix A.1), which is defined as the number of binary operations, $ACE=MACs\times bit_{weight}\times bit_{act.}$, where $bit_{weight}$ and $bit_{act.}$ indicates the bit-width of weights and activations respectively. (ii) For inference on neuromorphic chips, we consider the event-driven sparsity in SNNs. We introduce a Sparse ACE, defined as $sparsity\times ACE$, which is evaluated after expanding GIF neurons back to 1-bit spiking.
@@
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 0.0688$\times$ 65.45 41.67 32.51 64.37 56.59 54.3 52.48
LLAMA-2-7B FP16 1$\times$ 76.99 53.58 40.61 71.07 72.96 67.25 63.74
OmniQuant W4A4 0.0625$\times$ 61.75 40.32 29.86 60.89 48.86 52.17 48.98
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 0.0688$\times$ 62.35 41.41 29.95 58.87 54.27 50.20 49.51
OmniQuant W2A8 0.0625$\times$ 50.71 27.74 25.17 38.26 26.16 50.36 36.40
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A8 0.075$\times$ 53.37 31.27 23.63 53.79 33.94 51.46 41.24
OmniQuant W2A16 0.125$\times$ 57.02 32.83 25.34 53.46 31.50 50.36 41.75
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 0.138$\times$ 65.67 41.88 28.41 60.46 49.87 52.80 49.85
LLAMA-2-13B FP16 1$\times$ 79.05 57.95 44.28 69.02 76.63 69.61 66.09
OmniQuant W4A4 0.0625$\times$ 68.01 46.09 32.51 62.87 57.71 53.91 53.52
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 0.0688$\times$ 66.0 45.71 33.19 64.07 61.38 54.54 54.15
OmniQuant W2A8 0.0625$\times$ 56.96 33.0 23.55 61.31 32.0 49.41 42.71
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A8 0.075$\times$ 64.69 41.46 28.16 62.29 50.5 52.49 49.93
OmniQuant W2A16 0.125$\times$ 62.24 40.15 29.52 62.45 49.76 53.12 49.54
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 0.138$\times$ 67.68 44.53 30.38 63.64 58.02 57.22 53.58
LLAMA-2-70B FP16 1$\times$ 80.85 59.72 47.95 76.7 80.85 77.03 70.52
OmniQuant W2A16 0.125$\times$ 68.44 45.16 32.59 63.12 58.32 52.25 53.31
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 0.138$\times$ 75.84 51.68 39.42 66.67 68.44 60.22 60.38
Table 4: Comparisons with QuaRot. We follow evaluation metrics of QuaRot and all the results are evaluated by lm-eval-v0.4.2 following recent works. Saliency indicates the ratio of salient channels, and we set spiking step, T’=2, in salient channels.
[…]
LLAMA-2-7B - FP16 78.84 74.54 46.33 77.74 75.97 69.22 70.44
QuaRot-RTN - W4A4 71.82 59.89 36.18 67.37 63.88 59.12 59.71
SpikeLLM-RTN 0.2 W4A4 72.47 62.29 36.01 69.48 64.74 59.43 60.74
QuaRot-GPTQ - W4A4 75.95 68.43 39.76 72.54 72.23 64.72 65.61
SpikeLLM-GPTQ 0.2 W4A4 77.37 70.88 41.81 73.00 72.42 65.11 66.76
LLAMA-2-13B - FP16 80.63 77.48 49.23 80.73 79.37 71.74 80.69
QuaRot-RTN - W4A4 74.86 69.19 41.98 72.54 70.35 64.72 65.61
SpikeLLM-RTN 0.2 W4A4 75.79 69.53 41.21 74.31 71.51 65.51 66.31
QuaRot-GPTQ - W4A4 77.91 72.18 46.16 78.41 75.55 68.82 69.84
SpikeLLM-GPTQ 0.3 W4A4 78.45 73.74 46.84 78.75 76.40 68.03 70.37
Figure 4: Ablation studies in LLAMA-2-7B. Average accuracy (not norm) is reported. (a) Comparison between SpikeLLM and Quantized-ANN with the same operations. (b) Ablations on spiking salient channels in activations and KV-Caches. (c) Ablations on spiking salient channels in weights.
5.1 Main Results
Comparison with General Quantization. As shown in Table 3 and 10, we compare SpikeLLM with the most general weight-activation quantization methods including SmoothQuant, LLM-QAT, and OmniQuant, which shows SpikeLLM improves significantly based on OmniQuant with a few additional spikes to enhance salient channels. When replacing activation and KV-cache quantization with spiking neurons, SpikeLLM-W4A4 LLAMA-7B improves 2.55% accuracy and reduces 11.01% perplexity. When replacing weight quantization, improvements is more dramatic: for the LLAMA-2-7B-W2A8 and W2A16, $\text{SpikeLLM}_{\texttt{T=4}}$ and $\text{SpikeLLM}_{\texttt{T=2}}$ exceed 4.84% and 8.10% respectively; 89.58% and 61.89% perplexity are reduced. Notice that, SpikeLLM can plug in almost quantized LLMs beyond OmniQuant, which indicates the potential higher performance equipped with stronger quantization pipeline. In Appendix A.5, we further explore more training pipelines, such as QLoRA (Dettmers et al., 2024) and efficient QAT (Chen et al., 2024) settings.
Comparison with QuaRot. Recent QuaRot (Ashkboos et al., 2024) achieves outlier-free quantization with the help of additional Hadamard transformations. After rotation, the activation become a roughly ball-shape Gaussian distribution. In SpikeLLM framework, we select the more salient channels given a Gaussian distribution, and allocate more steps. As shown in Table 4, SpikeLLM improves 1.03% and 1.15% when quantizing weights by RTN and GPTQ respectively in LLAMA-2-7B, indicating the ability to cooperate with rotation-based quantization.
Table 5: Operations of SpikeLLM. ACE and Sparse ACE evaluate operations on the GEMM kernel and event-driven neuromorphic chips.
Models ACE ($10^{12}$) Sparse ACE ($10^{12}$)
LLAMA-1-7B 3886.22 3886.22
$\text{SpikeLLM}_{\texttt{T=2}}$ 360.20 608.78
$\text{SpikeLLM}_{\texttt{T=4}}$ 386.70 648.09
Operation Efficiency Compared with ANNs. Compared with LLAMA-2-7B, $\text{SpikeLLM}_{\texttt{T=2}}$ and $\text{SpikeLLM}_{\texttt{T=4}}$ for the W4A4 setting save ×$\times 6.38$ and ×$\times 10.79$ ACE in each layer. In Table 5, we evaluate the multi-bit and 1-bit workflows proposed in Sec. 4.1. Because GIF neurons expand much longer in 1-bit inference, it achieves asynchronous computation for the following linear layer at the cost of more operations than GEMM kernels.
Ablation Studies on GIF neurons. We evaluate the efficiency of GIF neurons by comparison with OmniQuant-W4A4 with almost the same operations. This is achieved by applying GIF neurons in both weights and activations. It is known that GIF neurons and OBSpiking always introduce a few additional operations. To confirm the same operations, we accordingly decrease the spike length in non-salient channels in weights and maintain the total spike length of weights and activations the same as OmniQuant-W4A4. As shown in Fig.4 (a), given different percentages of salient channels in activations, SpikeLLM always exceeds the Omniquant-W4A4 baseline shown as the red line.
Ablation Studies on Optimal Brain Spiking. To evaluate the efficiency of the OBSpiking framework, we compare SpikeLLM with equal-operation baselines with randomly selected spiking channels, termed Random-Spike. In Fig. 4 (b), we compare three settings: (i) SpikeLLM-W4A4: OBSpiking in activations of linear layers alone, (ii) SpikeLLM-KV-W4A4: OBSpiking in both activations of linear layers and KV caches of self-attentions, and (iii) Random-Spike in activations of linear layers alone. We don’t evaluate the random-spike in KV-caches because this setting is unstable to optimize. With different percentages of salient channels, OBSpiking in both KV chches and activations helps to improve performance. In Fig. 4 (c), we compare Random-Spike and OBSpiking in weight quantization, which proves the effectiveness for weights similarly.